# TNO, The Hague, The Netherlands

### v1

@@ 4 Conclusions and Recommendations
The total overview of all the literature discussed in the previous section can be found in Table 1 and Table 2. Here the paradigm, the main topic and the discussed problem are listed. Also, in these tables is indicated whether the solution is a pure quantum solution or a hybrid quantum solution. With hybrid we mean ‘Horizontal Hybrid Quantum Computing’, following the classification of phillipson2023classification . This means that the algorithm requires both a quantum computer and a classical computer to perform an algorithm, in sequential or parallel order. As we can see in the tables, most of the current solutions use hybrid approaches. This is as expected, from two thoughts. First, the current generation of quantum computers are still small, meaning that solving real world use cases, which is what most articles in the literature aim for, is limited and classical computers are needed to perform at least part of the algorithm. Next, for many optimisation problems a hybrid approach, using quantum computers as part of a (meta-) heuristic approach is expected to last, where also classical algorithms break down the calculation pipeline and the QC is not expected to solve NP-hard problems aaronson2008limits . These meta-heuristic approaches, where a QC is combined with a classical counterpart report the best performances. An other aspect that is clearly visible in this overview is that the majority of publications use QA as main quantum paradigm. This follows by the relative high number of qubits and larger size of problems this technology allows for. Also the hybrid approaches, as offered by D-Wave in their full stack approach, help with the uptake here. The QAOA and other VQE approaches on the gate-based device are the second most used approach, however they are still very much limited in performance. More problem specific and machine learning approaches are still limited in number. There is quite a bit of room for experimentation here. When we look at the topics and problems covered, we see most attention to the classical areas of routing and scheduling. Especially in predictions there is again room for further research, potentially for QML approaches.
Generally, we can say that we endorse the conclusions reached in the work of Osaba osaba2022systematic . Were we looked more general than only VRP and TSP work, also we can conclude that the majority of the papers conclude that current hardware is in the early stage: researchers face small number of qubits and inherent limitations such as noise and decoherence, and challenges in tailoring problems to hardware capacities. This leads to limitations in addressing complex formulations. Only a small number of papers indicate that their hybrid approach beats current classical approaches. Most of the work indicate the potential advantages and benefits of their approach.
Acknowlegements
@@
Ajagekar ajagekar2020quantum2 QA - QUBO Routing VRP hybrid
Alsaiyari alsaiyari2023variational GBC - QAOA/VQE Routing VRP hybrid
Atchade-Adelomou atchade2021qrobot QA and GBC - QUBO Routing picking hybrid
Azad azad2022solving GBC - QAOA Routing VRP hybrid
Azzaoui azzaoui2021quantum GBC - QAOA Routing TSP hybrid
Bentleybentley2022quantum GBC - QAOA Routing CVRP hybrid
Borowski borowski2020new QA - QUBO Routing VRP/CVRP full and hybrid
Bourreau bourreau2023indirect GBC - QAOA Routing TSP hybrid
Correll correll2023quantum GBC - QML Routing VRP hybrid
Dixit dixit2023quantum2 QA - QUBO Routing STDSP full
Feld feld2019hybrid QA - QUBO Routing CVRP hybrid
Fitzek fitzek2021applying GBC - QAOA Routing HVRP hybrid
Harikrishnakumar harikrishnakumar2020quantum QA - QUBO Routing (D-)MDCVRP full
Harwood harwood2021formulating GBC - QAOA Routing VRPTW hybrid
Irie irie2019quantum QA - QUBO Routing CVRPTW hybrid
Lo lo2021genetic QRNG Routing VRP hybrid
Khumalo khumalo2022investigation GBC - QAOA/VQE Routing TSP hybrid
Le le2023quantum QA - QUBO Routing sTSP hybrid
Leonidas leonidasqubit GBC - QAOA Routing VRPTW hybrid
Lo lo2021genetic QRNG Routing VRP hybrid
Makhanov makhanov2023quantum GBC - Grover’s Routing SPP hybrid
Marsoit marsoit2021quantum QA - QUBO Routing CVRP hybrid
Masuda masuda85optimisation CA - QUBO Routing TDVRPTW full
Mohanty mohanty2023analysis GBC VQE Routing VRP hybrid
Neukart neukart2017traffic QA - QUBO Routing Flow opt. hybrid
Palackal palackal2023quantum GBC - QAOA/VQE Routing TSP hybrid
Palmieri palmieriquantum QA and GBC - QUBO Routing CVRP hybrid
Papalitsas papalitsas2019qubo QA - QUBO Routing TSPTW full
Phillipson phillipson2021multimodal QA - QUBO Routing Flow opt. hybrid
Poggel poggel2023recommending Generic Routing CVRP hybrid
Rana rana2022quantum QA and GBC - QUBO Routing CVRP hybrid
Salehi salehi2022unconstrained QA - QUBO Routing TSPTW full
Sales sales2023adiabatic QA - QUBO Routing CVRP full and hybrid
Sanches sanches2022short GBC - RL Routing CVRP hybrid
Sinno sinno2023performance QA - QUBO Routing CVRP hybrid
Spyridis spyridis2023variational GBC - QAOA Routing m-TSP hybrid
Srinivasan srinivasan2018efficient GBC - QPE Routing TSP hybrid
Warren warren2020solving QA - QUBO Routing TSP full and hybrid
Weinberg weinberg2023supply QA - QUBO Routing CVRP hybrid
Xie xie2023feasibility GBC - QAOA Routing CVRP hybrid
Yarkoni yarkoni2021solving QA - QUBO Routing SRP hybrid
Choo chooinvestigating QA - QUBO Network Design FLP hybrid
Ding ding2021implementation QA -QUBO Network Design NDP hybrid
Ding ding2019logistic QA -QUBO Network Design NDP full
Dixit dixit2023quantum QA -QUBO Network Design NDP hybrid
Gabbassov gabbassov2022transit QA - QUBO Network Design UTP hybrid
Giraldo giraldo2022using QA and GBC - QUBO Network Design MCLP hybrid
Guo guo2020quantum QA - QUBO Network Design FLP hybrid
Khumalo khumalo2022investigation GBC - QAOA/VQE Network Design FLP hybrid
Klar klar2022quantum QA - QUBO Network Design FLP hybrid
Mahasinghe mahasinghe2021qubo QUBO Network Design FLP -
Malviya malviya2023logistics QA - QUBO Network Design NDP hybrid
Satori satori2023quantum QA - QUBO Network Design LAP hybrid
Tosum tosun2022new QUBO Network Design QAP -
@@
Martins martins2020applying QA - QUBO Fleet opt. TAP hybrid
Vikstaal vikstaal2020applying GBC - QAOA Fleet opt. TAP hybrid
Willsch willsch2022benchmarking QA - QUBO Fleet opt. TAP hybrid
De Andoin de2022comparative QA - QUBO Cargo BPP hybrid
De Andoin de2022hybrid QA - QUBO Cargo BPP hybrid
Benson benson2023cqm QA - CQM Cargo KP hybrid
Bontekoe bontekoe2023translating QA - QUBO Cargo QKP full
Cellini cellini2023qal QA - QUBO Cargo BPP full
Nayak nayak2022quantum QA - QUBO Cargo ALO hybrid
Pilon pilon2021aircraft QA - QUBO Cargo ALO full
Romero romero2023solving ; v2023hybrid QA - CQM Cargo BPP hybrid
Sotelo hernandez2020application ; sotelo2021determination GBC - VQE Cargo ALO hybrid
Van Dam van2021quantum GBC - QAOA Cargo KP hybrid
Jahin jahin2023qamplifynet GBC - QML Prediction Prediction hybrid
Jiang jiang2022quantum GBC - HHL/VQE Prediction Inventory hybrid
Adelomou adelomou2020using GBC - VQE Scheduling Personnel hybrid
Ajagekar ajagekar2020quantum2 QA - QUBO Scheduling JSP hybrid
Ajagekar ajagekar2020quantum2 QA - QUBO Scheduling MCFP hybrid
Amaro amaro2022case GBC - VQE Scheduling JSP hybrid
Carugno carugno2022evaluating QA - QUBO Scheduling JSP hybrid
Denkena denkena2021quantum QA - QUBO Scheduling FJSP hybrid
Huang huang2022paint QA - QUBO Scheduling JSP hybrid
Kurowski kurowski2020hybrid QA - QUBO Scheduling JSP hybrid
Mohammadbagherpoor mohammadbagherpoor2021exploring QA - VQE Scheduling QAP hybrid
Mori mori2022replanning GBC and QA - QAOA Scheduling ARP hybrid
Permin permin2022simple GBC - Grover Scheduling JSP hybrid
Riandari riandari2021quantum GBC - QAOA Scheduling Production opt. hybrid
Scherer scherer2021oncall GBC - Grover Scheduling JSP full
Streif streif2021beating GBC - QAOA Scheduling JSP hybrid
Schworm schworm2023solving QA - QUBO Scheduling JSP hybrid
Venturelli venturelli2015quantum QA - QUBO Scheduling JSP full
Yarkoni yarkoni2021multi QA - QUBO Scheduling JSP hybrid

### v2

@@ 4 Conclusions and Recommendations
The total overview of all the literature discussed in the previous section can be found in Table 1 and Table 2. Here the paradigm, the main topic and the discussed problem are listed. Also, in these tables is indicated whether the solution is a pure quantum solution or a hybrid quantum solution. With hybrid we mean ‘Horizontal Hybrid Quantum Computing’, following the classification of [114]. This means that the algorithm requires both a quantum computer and a classical computer to perform an algorithm, in sequential or parallel order. As we can see in the tables, most of the current solutions use hybrid approaches. This is as expected, from two thoughts. First, the current generation of quantum computers are still small, meaning that solving real world use cases, which is what most articles in the literature aim for, is limited and classical computers are needed to perform at least part of the algorithm. Next, for many optimisation problems a hybrid approach, using quantum computers as part of a (meta-) heuristic approach is expected to last, where also classical algorithms break down the calculation pipeline and the QC is not expected to solve NP-hard problems [1]. These meta-heuristic approaches, where a QC is combined with a classical counterpart report the best performances. An other aspect that is clearly visible in this overview is that the majority of publications use QA as main quantum paradigm. This follows by the relative high number of qubits and larger size of problems this technology allows for. Also the hybrid approaches, as offered by D-Wave in their full stack approach, help with the uptake here. The QAOA and other VQE approaches on the gate-based device are the second most used approach, however they are still very much limited in performance. More problem specific and machine learning approaches are still limited in number. There is quite a bit of room for experimentation here. When we look at the topics and problems covered, we see most attention to the classical areas of routing and scheduling. Especially in predictions there is again room for further research, potentially for QML approaches.
Generally, we can say that we endorse the conclusions reached in the work of Osaba et al. [107]. Where we looked more general than only VRP and TSP work, also we can conclude that the majority of the papers conclude that current hardware is in the early stage: researchers face small number of qubits and inherent limitations such as noise and decoherence, and challenges in tailoring problems to hardware capacities. This leads to limitations in addressing complex formulations. Only a small number of papers indicate that their hybrid approach beats current classical approaches. Most of the work indicate the potential advantages and benefits of their approach.
Acknowledgements
@@
Ajagekar [4] QA - QUBO Routing VRP hybrid
Alsaiyari [6] GBC - QAOA/VQE Routing VRP hybrid
Atchade-Adelomou [11] QA and GBC - QUBO Routing picking hybrid
Azad [12] GBC - QAOA Routing VRP hybrid
Azzaoui [13] GBC - QAOA Routing TSP hybrid
Bentley[16] GBC - QAOA Routing CVRP hybrid
Borowski [19] QA - QUBO Routing VRP/CVRP full and hybrid
Bourreau [20] GBC - QAOA Routing TSP hybrid
Cattelan [29] QA - QUBO Routing RPP hybrid
Correll [36] GBC - QML Routing VRP hybrid
Dixit [46] QA - QUBO Routing STDSP full
Feld [49] QA - QUBO Routing CVRP hybrid
Fitzek [50] GBC - QAOA Routing HVRP hybrid
Fitzek [51] GBC - QAOA Routing HVRP hybrid
Harikrishnakumar [62] QA - QUBO Routing (D-)MDCVRP full
Herzog [65] GBC - QAOA Routing CVRP hybrid
Harwood [63] GBC - QAOA Routing VRPTW hybrid
Irie [67] QA - QUBO Routing CVRPTW hybrid
Kanai [73] QA - QUBO Routing CVRP hybrid
Khan [74] GBC -Grover’s/QAOA Routing FRO hybrid
Khumalo [75] GBC - QAOA/VQE Routing TSP hybrid
Le [81] QA - QUBO Routing sTSP hybrid
Leonidas [82] GBC - QAOA Routing VRPTW hybrid
Li [83] GBC - VQA Routing VRP hybrid
Liu [84] QA - QUBO Routing TSP hybrid
Lo [85] QRNG Routing VRP hybrid
Makhanov [87] GBC - Grover’s Routing SPP hybrid
Makhanov [88] GBC - Grover’s Routing FRO hybrid
Mario [91] QA - QUBO Routing CVRP hybrid
Marsoit [92] QA - QUBO Routing CVRP hybrid
Masuda [95] CA - QUBO Routing TDVRPTW full
Mohanty [99] GBC - VQE Routing VRP hybrid
Mohanty [100] GBC - QSVM Routing VRP hybrid
Neukart [104] QA - QUBO Routing Flow opt. hybrid
Osaba [106] QA - QUBO Routing CVRP hybrid
Palackal [108] GBC - QAOA/VQE Routing TSP hybrid
Palmieri [109] QA and GBC - QUBO Routing CVRP hybrid
Papalitsas [110] QA - QUBO Routing TSPTW full
Phillipson [113] QA - QUBO Routing Flow opt. hybrid
Poggel [117] Generic Routing CVRP hybrid
Qiu [118] GBC - QACO Routing TSP hybrid
Ramezani [120] GBC - QAOA Routing TSP hybrid
Rana [121] QA and GBC - QUBO Routing CVRP hybrid
Rosendo [125] QA and GBC - QUBO Routing CVRP hybrid
Sadashivan [126] QA - QUBO Routing CVRP hybrid
Salehi [128] QA - QUBO Routing TSPTW full
Sales [129] QA - QUBO Routing CVRP full and hybrid
@@
Salloum [130] QA - QUBO Routing TFO hybrid
Salloum [131] QA - QUBO Routing TFO hybrid
Sanches [132] GBC - RL Routing CVRP hybrid
Sato [133] GBC - Quantum Search Routing TSP full
Sinno [144] QA - QUBO Routing CVRP hybrid
Spyridis [146] GBC - QAOA Routing m-TSP hybrid
Srinivasan [147] GBC - QPE Routing TSP hybrid
Tejani [151] GBC - QPE Routing BTSP hybrid
Warren [158] QA - QUBO Routing TSP full and hybrid
Weinberg [159] QA - QUBO Routing CVRP hybrid
Wesolowski [160] GBC Routing SPSP full
Xie [163] GBC - QAOA Routing CVRP hybrid
Xu [166] GBC - QML Routing CVRP hybrid
Yarkoni [168] QA - QUBO Routing SRP hybrid
Buß [25] GBC - QAOA Network Design FLP hybrid
Chiscop [33] QA - QUBO Network Design MSLSCP hybrid
Choo [34] QA - QUBO Network Design FLP hybrid
Ding [44] QA -QUBO Network Design NDP hybrid
Ding [43] QA -QUBO Network Design NDP full
Dixit [45] QA -QUBO Network Design NDP hybrid
Gabbassov [52] QA - QUBO Network Design UTP hybrid
Giraldo [54] QA and GBC - QUBO Network Design MCLP hybrid
Guo [59] QA - QUBO Network Design FLP hybrid
Khumalo [75] GBC - QAOA/VQE Network Design FLP hybrid
Klar [76] QA - QUBO Network Design FLP hybrid
Mahasinghe [86] QA - QUBO Network Design FLP -
Malviya [90] QA - QUBO Network Design NDP hybrid
Malviya [89] GBC - QAOA Network Design NDP full
Radvand [119] GBC -Grover’s Network Design CSLP full
Sakib [127] QA - QUBO Network Design CSLP hybrid
Satori [134] QA - QUBO Network Design LAP hybrid
Tosum [153] QA - QUBO Network Design QAP -
Wang [157] GBC - QAOA Network Design FLP hybrid
Martins [94] QA - QUBO Fleet opt. TAP hybrid
Vikstaal [156] GBC - QAOA Fleet opt. TAP hybrid
Willsch [161] QA - QUBO Fleet opt. TAP hybrid
De Andoin [8] QA - QUBO Cargo BPP hybrid
De Andoin [9] QA - QUBO Cargo BPP hybrid
Ardelian [10] GBC - QML Cargo QKP hybrid
Benson [15] QA - CQM Cargo KP hybrid
Bontekoe [18] QA - QUBO Cargo QKP full
Bozejko [21] QA - QUBO Cargo QKP hybrid
Cellini [30] QA - QUBO Cargo BPP full
Cellini [31] QA - QUBO Cargo BPP hybrid
Christiansen [35] GBC - QAOA Cargo QKP hybrid
Cui [38] GBC - Grover’s Cargo QKP hybrid
Gatti [53] QA - QUBO Cargo BPP hybrid
@@
Work Paradigm Topic Problem Nature
Matt [96] QAOA Cargo BPP hybrid
Nayak [103] QA - QUBO Cargo ALO hybrid
Pilon [116] QA - QUBO Cargo ALO full
Romero [123, 124] QA - CQM Cargo BPP hybrid
Shirai [141] GBC - VQA Cargo QKP hybrid
Sotelo [64, 145] GBC - VQE Cargo ALO hybrid
Van Dam [40] GBC - QAOA Cargo KP hybrid
Gutta [60] GBC - QML Prediction Forecasting hybrid
Jahin [69] GBC - QML Prediction Prediction hybrid
Jiang [71] GBC - HHL/VQE Prediction Inventory hybrid
Koushik [78] GBC - QML Prediction Maintenance hybrid
Sehrawat [140] GBC - QML Prediction Forecasting hybrid
Adelomou [2] GBC - VQE Scheduling Personnel hybrid
Ajagekar [4] QA - QUBO Scheduling JSP hybrid
Ajagekar [4] QA - QUBO Scheduling MCFP hybrid
Amaro [7] GBC - VQE Scheduling JSP hybrid
Bernreuther [17] GBC - QAOA Scheduling Scheduling hybrid
Carugno [27] QA - QUBO Scheduling JSP hybrid
Denkena [42] QA - QUBO Scheduling FJSP hybrid
Grange [57] GBC - QAOA Scheduling Scheduling hybrid
Huang [66] QA - QUBO Scheduling JSP hybrid
Krol [79] GBC - Grover’s Scheduling Scheduling full
Kurowski [80] QA - QUBO Scheduling JSP hybrid
Mohammadbagherpoor [98] QA - VQE Scheduling QAP hybrid
Mori [101] GBC and QA - QAOA Scheduling ARP hybrid
Permin [111] GBC - Grover Scheduling JSP hybrid
Riandari [122] GBC - QAOA Scheduling Production opt. hybrid
Scherer [135] GBC - Grover Scheduling JSP full
Streif [148] GBC - QAOA Scheduling JSP hybrid
Schworm [137] QA - QUBO Scheduling JSP hybrid
Schworm [138] QA - QUBO Scheduling JSP hybrid
Schworm [139] QA - QUBO Scheduling JSP hybrid
Venturelli [155] QA - QUBO Scheduling JSP full
Windmann [162] GBC - QAOA Scheduling JSP hybrid
Yarkoni [167] QA - QUBO Scheduling JSP hybrid
Table 3: Overview - Part III
@@
MDKP Multi-Dimensional Knapsack Problem
MSLSCP Multi-Service Location Set Covering Problem